OpenStudy (anonymous):

Find the zeros of the function. State the multiplicity of multiple zeros. y = (x - 4)2 I found the zeros, (x =4) but I'm not sure how to find the multiplicity

ganeshie8 (ganeshie8):

(x-4)^2 = (x-4)(x-4) = 0 x = 4, 4

ganeshie8 (ganeshie8):

4 repeated "2" times, so the multiplicity is "2"

OpenStudy (anonymous):

Oh! I didn't know it was that easy... Thank you!

ganeshie8 (ganeshie8):

yw!